The length of the base of a triangle is twice its height. If the area is 100 square kilometers, find the height.
Found 2 solutions by unlockmath, CharlesG2:
Answer by unlockmath(1688)

Hello,
Since we know the Area of a triangle is equal to 1/2(Base times height) we can plug in the values known and let "x" represent the height:
100=1/2(x)(2x) Rewritten as:
100=x^2
Square root each side to get:
x=10
So height is 10 km
Width is 20 km
Make sense?
There's a book I wrote that might help;
www.math-unlock.com

Answer by CharlesG2(834)

The length of the base of a triangle is twice its height. If the area is 100 square kilometers, find the height.
A area = 1/2 * B base * H height
100 = 1/2 * (2H) * H
100 = H^2
10 = H
height is 10 km